OpenStudy (anonymous):

2= n+2/ 6 I dont get itt.

OpenStudy (anonymous):

multiply 6 and 2 12=n+2 subtract 2 on both sides n=10

OpenStudy (anonymous):

\[2=n+\frac{2}6\] \[2-\frac{2}6=n\] \[\frac{12-2}6=n\] \[\frac{10}6=n\] \[\frac{5}{3}=n\]

OpenStudy (anonymous):

is it (n+2)/6 ??

OpenStudy (anonymous):

and if u meant this : \[2=\frac{n+2}6\] \[2*6=n+2\] \[12=n+2\] \[12-2=n\] \[10=n\]

OpenStudy (anonymous):

multiply 6 times 2 which equals 12 12=n+2 then subtract 2 from 12 which equals 10=n